# beSnake: A routing algorithm for scalable spin-qubit architectures

### v1

@@ VI Conclusion
This paper presents beSnake, a novel routing algorithm designed explicitly for scalable spin-qubit architectures. beSnake addresses the unique challenges posed by spin-qubit systems, such as the incorporation of the shuttle operation over the traditional SWAP gate used in other qubit technologies. We presented beSnake’s ability to dynamically adapt to different routing challenges and efficiently handle complex scenarios involving multiple parallel gates. We stress-tested beSnake in various configuration options to provide valuable performance insights. In particular, we showed that adjusting the heuristic of the shortest path selection can offer a significant speed improvement with a relatively small overhead cost. Then, beSnake was tested on its SWAP replacement option and its capacity to handle a highly connected topology faster than a less connected one. Furthermore, we gave an insight into the optimal qubit density based on all metrics that beSnake can route in a scalable manner. In the second simulation phase, we conducted an extensive comparison between the Shuttle-based SWAP (SBS) routing algorithm acting on random and real algorithms with up to 1000 qubits, in which beSnake demonstrated significant improvements, with up to and on average in gate overhead and depth overhead, respectively, and up to times faster routing time. We plan to make beSnake publicly available and integrated into SpinQ [1].
While performing our research, several avenues for future work have emerged, which warrant further exploration and investigation. Regarding the mechanisms for dealing with blockages discussed in IV.4.1, the list of alternative approaches can be further expanded for such edge cases. This should be done with scalability in mind in terms of the extra computation time and whether it will be worth it based on their utility on the particular topological patterns and size of the coupling graph. One example is to move qubits for two or more gates concurrently, or to allow the operand qubits to be pushed by other qubits. This will possibly improve the output circuit’s overhead but also make it slower for beSnake to find a solution. In other cases, it can increase both overheads because of path conflicts and deeper BFS explorations.
As for ways to sort parallelized Z and two-qubit gates, besides the one discussed in IV.4.2, more nuanced techniques will most likely increase the algorithm’s complexity but improve both overheads. One possible direction is to create a heuristic upon which the order of all the different gate types is sorted based on their predicted routing paths. Regardless, such strategies and those discussed previously can be explored and compared in targeted architectures that will use them.

### v2

@@ VI Conclusion
This paper presents beSnake, a novel routing algorithm designed explicitly for scalable spin-qubit architectures. beSnake addresses the unique challenges posed by spin-qubit systems, such as the incorporation of the shuttle operation over the traditional SWAP gate used in other qubit technologies. We presented beSnake’s ability to dynamically adapt to different routing challenges and efficiently handle complex scenarios involving multiple parallel gates. We stress-tested beSnake in various configuration options to provide valuable performance insights. In particular, we showed that adjusting the heuristic of the shortest path selection can offer a significant speed improvement with a relatively small overhead cost. Then, beSnake was tested on its SWAP replacement option and its capacity to handle a highly connected topology faster than a less connected one. Furthermore, we gave an insight into the optimal qubit density, up to , that beSnake can route in a scalable manner. In the second simulation phase, we conducted an extensive comparison between the SBS routing algorithm acting on random and real algorithms with up to 1000 qubits, in which beSnake demonstrated significant improvements, with up to and on average in gate overhead and depth overhead, respectively, and up to times faster routing time.
While performing our research, several avenues for future work have emerged, which warrant further exploration and investigation. Regarding the mechanisms for dealing with blockages discussed in IV.4.1, the list of alternative approaches can be further expanded for such edge cases. This should be done with scalability in mind in terms of the extra computation time and whether it will be worth it based on their utility on the particular topological patterns and size of the coupling graph. One example is to move qubits for two or more gates concurrently, or to allow the operand qubits to be pushed by other qubits. This will possibly improve the output circuit’s overhead but also make it slower for beSnake to find a solution. In other cases, it can increase both overheads because of deeper BFS explorations due to path conflicts within the different layers.
As for ways to sort parallelized Z and two-qubit gates contained in a cycle, besides the one discussed in IV.4.2, more nuanced techniques will most likely increase the algorithm’s complexity but improve both overheads. One possible direction is to create a heuristic upon which the order of all the different gate types is sorted based on their predicted routing paths. Regardless, such strategies and those discussed previously should be explored and compared in targeted architectures that will use them.